OpenStudy (abb0t):

connexus algebra 2b help plez!!!

OpenStudy (abb0t):

\[Simplify:\large \frac{ (4x^2+1)^{\frac{ 1 }{ 2 }}(6x)-3x^2(\frac{ 1 }{ 2 })(4x^2+1)^{-\frac{ 1 }{ 2 }}(8x) }{ \left[ (4x^2+1)^\frac{ 1 }{ 2 } \right]^2 }\]

OpenStudy (anonymous):

Are you mocking all those connexus students in Algebra II B? Because I know for a fact that isn't one of the questions.

OpenStudy (anonymous):

....lol

OpenStudy (primeralph):

for ril?